Reports 1-1 of 1 Clear search Modify search
Detector Characterisation (Broadband noise)
ruggi - 13:25 Saturday 10 August 2019 (46638) Print this report
Drop of sensitivity

The worsening of the sensitivity, centered at 100 Hz (fig 2), cannot be found at all in DARM (fig 3). The drop of the optical gain (fig 4) has to be taken in account, but probably this is not enough to explain the discrepancy. Given that Hrec is the result of a noise subtraction, a different behavior with respect to DARM is expected, but in any case it could be worth to give a deeper look.

A similar worsening is not present in PRCL and MICH (fig 5, fig 6), as one could expect in case the extra noise in Hrec was related to some excess of scattered light after the modification of input power.

Images attached to this report
Comments to this report:
mwas - 15:42 Saturday 10 August 2019 (46640) Print this report

Figure 1. In addition to the optical gain going down by 5% the power in the arms increased by 5%. Normally an increase in power in the arms should lead to an increase in the optical gain, so the drop in optical gain is larger than it seems. The DARM UGF drop by 5% too, most likely for the same reasons as the optical gain drop.


Images attached to this comment
tacca, nardecchia - 10:19 Sunday 11 August 2019 (46644) Print this report
We made two different steps with the rotator to restore the input power: in the first step the power has been increased from about 18.5 W to about 20 W (entry 46616), while in the second step the power has been decreased to about 19.5 W (entry 46631). Looking at Michal's plot, after the first step (input power = 20 W) also the optical gain increased.
In the attached plots there are hrec and DARM in three different cases:
- Blue curve: Aug 6th, 18:30 UTC -> 48.5 Mpc;
- Red curve: Aug 8th, 21:30 UTC -> 45 Mpc;
- Yellow curve: Aug 10th, 7:30 UTC -> 42 Mpc.
The CMRF is good in the current sensitivity (suggesting that the ITF is in a good working point from ISC and TCS point of view), while it was spoiled on Aug 8th (one of the reasons that pushed us to decrease a bit the input power, with a general instability of the ITF).
As Paolo already pointed out, between Aug 6th and Aug 10th, DARM is not really different above 60 Hz, while on Aug 8th DARM seemed to be much worse.
Images attached to this comment
ruggi - 19:50 Wednesday 14 August 2019 (46671) Print this report

The excess of noise in Hrec at 100-150 Hz follows the trend of PR to h coupling (fig 1). It was not that with the old version of Hrec (and this is expected, because PRCL is no more subtracted in that frequency range). This can be seen also comparing hoft to a roughly calibrated DARM (fig 2), and computing transfer function and coherence from PRCL to hoft and PRCL to DARM respectively (fig 3).

Images attached to this comment
mwas - 15:41 Thursday 15 August 2019 (46676) Print this report

Indeed the contribution from PRCL has increased, and is seems to be back at the level it was at the beginning of the run.

Figure 1. Shows an offline noise subtraction where channels are considered at all frequencies the inverse of the matrix of transfer function between all the noise subtractions channels is used to be use correlated channels simultaneously (which the online Hrec subtraction is not able to do at the present).  The contribution of PRCL (green) becomes larger than the contribution of MICH at 90Hz, instead of at 150Hz as it was 2 weeks ago.

Figure 2, Shows the ration of h(t) cleaned online and offline in red, the offline cleaning improves the range by about 2Mpc.

This can be compared to figure 3 and 4, which are the same as figure 1 and 2 but for Aug 5 when the range was at ~48Mpc. So better noise subtraction could bring us from ~42Mpc, to ~44Mpc, but there is still 3-4Mpc missing.

Figure 5, compares the offline cleaing between Aug 11(red) and Aug 5 (black). There is clear higher shot noise now, which should be due to poorer optical gain, but this is not explaining the remaining difference in the 80Hz-200Hz bucket.

In any case, switching back the cross-over between SPRB_B4_56Mhz_Q (MICH) and SIB2_B2_8MHz_I (PRCL) subtraction in h(t) from 150Hz back to 90Hz where it was since the beginning of the run, should improve the range by ~1Mpc.


Images attached to this comment
tacca - 13:29 Saturday 17 August 2019 (46689) Print this report
In the lock after the troubles of the automation of tonight and the restart of all the VPM metatron processes, the ITF working point probably changed. The optical gain increased, the PR coupling decreased and the BNS range increased by a couple of Mpc. To investigate better.
Images attached to this comment
ruggi - 13:51 Saturday 17 August 2019 (46690) Print this report

At about 7.30 local time the IMC transmitted power increased a bit

Images attached to this comment
cleva, Derossi - 19:03 Tuesday 20 August 2019 (46724) Print this report

it seems to be connected to some "sudden" misalignement of the seeder wrt the neoVAN which, at least for this time, goes into the right direction

PMC_TRA increased together with PMC_REFL decreased means either the matching or the aligment have changed. (plot1)

A zoom (plot2) shows that an increase of the neoVAN power (no known reason so far) seems to have triggered the beam drift which has ended up in a better alignement of the PMC

Rem.: the time Cst of the beam drift is ~5 mn, while the time Cst noticed in 46722  (neoVAN beam drift vs pump diode power calibration attempt) is few seconds


Images attached to this comment
cleva - 18:25 Tuesday 03 September 2019 (46831) Print this report

plot3 points out the lower correction provided by the PSTAB after the 17/08 7h30 UTC.

It is connected to the better alignement of the PMC already noticed (see 46690  and coment).

Plot 1 & 2, before and after the power increase, (alignment improvement) witness the situation improvement

Images attached to this comment
robinet - 9:15 Monday 09 September 2019 (46864) Print this report

Indeed the situation improved nicely... until the recent bad weather conditions. High frequency glitches (+BNS range drops) appeared again over the data of this week-end. They are likely due to PMC mis-alignments and large PSTAB corrections.

cleva - 17:58 Monday 16 September 2019 (46933) Print this report

for sure the PMC alignement should be optimized.

we see the instability of the power raising along with the decrease/increase of PMC_T/PMC_R


Images attached to this comment
Search Help